Silver Sun Recycling has a weighted-average cost of capital of 8.34 percent and is evaluating two projects: A and B. Project A involves an initial investment of 4,818 dollars and an expected cash flow of 8,865 dollars in 6 years. Project A is considered more risky than an average-risk project at Silver Sun Recycling, such that the appropriate discount rate for it is 1.22 percentage points different than the discount rate used for an average-risk project at Silver Sun Recycling. The internal rate of return for project A is 10.7 percent. Project B involves an initial investment of 5,885 dollars and an expected cash flow of 8,710 dollars in 5 years. Project B is considered less risky than an average-risk project at Silver Sun Recycling, such that the appropriate discount rate for it is 2.18 percentage points different than the discount rate used for an average-risk project at Silver Sun Recycling. The internal rate of return for project B is 8.16 percent. What is X if X equals the NPV of project A plus the NPV of project B?
Answer #1

Discount rate for project A = WACC + 1.22% (Given Project A is more risky so we have to add)

= 8.34% + 1.22%

= 9.56%

NPV for Project A = Present value of future cash flows - initial cash outflow

Present value = future value / (1+r)^n

where , r = Discount rate

n = number of periods

= 8865 / (1+9.56%)^6 - 4818

= 5125.86 - 4818

= 307.86

Disccount rate for project B = 8.34% - 2.18% = 6.16%

NPV = 8710 / (1+6.16%)^5 - 5885

= $574.72

so X = 307.86 + 574.72

= $882.58 (rounded to two decimals)